OpenStudy (shaw74):

What is the weight (in grams) of a liquid that exactly fills a 182.8 milliliter container if the density of the liquid is 0.135grams over milliliter? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (shaw74):

i know density equals mass divided by volume

OpenStudy (anonymous):

Density = (Mass)/(Volume) 0.135 = M/182.8 solve for M

OpenStudy (anonymous):

DO you understand now?

OpenStudy (shaw74):

24.678

OpenStudy (shaw74):

how do i round if the answer is correct

OpenStudy (anonymous):

24.7

OpenStudy (anonymous):

medal? lol

OpenStudy (aum):

"Round to the nearest hundredth" means two decimals 24.678 rounds to 24.68 grams.

OpenStudy (shaw74):

can you check one more

OpenStudy (anonymous):

sure make a new question though

OpenStudy (shaw74):

A large emerald with a mass of 378.24 grams was recently discovered in a mine. If the density of the emerald is 2.76grams over centimeters cubed, what is the volume? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (anonymous):

close this one

OpenStudy (shaw74):

is it 137.04

OpenStudy (anonymous):

make a new question

OpenStudy (anonymous):

no what did you put for mass and density

OpenStudy (anonymous):

the volume is not given so you have to input the mass and density

OpenStudy (shaw74):

yes i divided mass and density

OpenStudy (anonymous):

then yes that is the correct answer

OpenStudy (shaw74):

how do i round it